# Square-free orders for CM elliptic curves modulo $p$ in short intervals

### v1

@@ 4. Proof of Theorem 2
~$\tilde{\mathcal{E}}(x,m^{2})=\pi_{D_{m^{2}}}(x+h)-\pi_{D_{m^{2}}}(x)-\frac{|D_{m^{2}}|(\operatorname{Li}(x+h)-\operatorname{Li}(x))}{[K(E[m^{2}]):K]}.$
As argued in Section 3, ~$\tilde{\mathcal{E}}(x,m^{2})$ is equal to
@@
(31) $\begin{split}\mathcal{M}(x,h,y,2\sqrt{x+h})&\leq\sum_{\begin{subarray}{c}y<m\leq 2\sqrt{x+h}\\ \text{$m$ square-free}\\ m=m_{i}m_{r}m_{s}\end{subarray}}\sum_{\begin{subarray}{c}x<p\leq x+h\\ a_{p}\neq 0\\ (\pi_{p}-1)=m_{i}m_{r}I(m_{s})\cdot(\alpha)\end{subarray}}1\\ &\ll\sum_{\begin{subarray}{c}y<m\leq 2\sqrt{x+h}\\ \text{$m$ square-free}\\ m=m_{i}m_{r}m_{s}\end{subarray}}3^{\nu(m_{s})}\Big{(}\frac{h}{m_{i}^{2}m_{r}^{2}m_{s}^{2}}+\frac{\sqrt{x}}{m_{i}m_{r}m_{s}}\Big{)}\\ &\ll\sum_{m_{s}\leq 2\sqrt{x+h}}\sum_{\frac{y}{m_{s}}<\tilde{m}\leq 2\sqrt{x+h}}3^{\nu(m_{s})}\Big{(}\frac{h}{\tilde{m}^{2}m_{s}^{2}}+\frac{\sqrt{x}}{\tilde{m}m_{s}}\Big{)}\\ &\ll\sum_{m_{s}\leq 2\sqrt{x+h}}3^{\nu(m_{s})}\Big{(}\frac{h}{ym_{s}}+\frac{\sqrt{x}}{m_{s}}\log x\Big{)}\\ &\ll\frac{h}{y}(\log x)^{3}+\sqrt{x}(\log x)^{4}.\end{split}$
Note that the last bound follows from the estimates $\sum_{m\leq x}\frac{1}{m}=\log x+O(1)$ and
@@ 5. Proof of Theorem 3
In [18], the author used the estimate
@@ Acknowledgments
The author would like to thank Professors Wen-Ching Winnie Li and Robert C. Vaughan for their suggestions.

### v2

@@ 4. Proof of Theorem 2
$\tilde{\mathcal{E}}(x,m^{2})=\pi_{D_{m^{2}}}(x+h)-\pi_{D_{m^{2}}}(x)-\frac{|D_{m^{2}}|(\operatorname{Li}(x+h)-\operatorname{Li}(x))}{[K(E[m^{2}]):K]}.$
As argued in Section 3, $\tilde{\mathcal{E}}(x,m^{2})$ is equal to
@@
(31) $\begin{split}\mathcal{M}(x,h,y,2\sqrt{x+h})&\leq\sum_{\begin{subarray}{c}y<m\leq 2\sqrt{x+h}\\ \text{$m$ square-free}\\ m=m_{i}m_{r}m_{s}\end{subarray}}\sum_{\begin{subarray}{c}x<p\leq x+h\\ a_{p}\neq 0\\ (\pi_{p}-1)=m_{i}m_{r}I(m_{s})\cdot(\alpha)\end{subarray}}1\\ &\ll\sum_{\begin{subarray}{c}y<m\leq 2\sqrt{x+h}\\ \text{$m$ square-free}\\ m=m_{i}m_{r}m_{s}\end{subarray}}3^{\nu(m_{s})}\Big(\frac{h}{m_{i}^{2}m_{r}^{2}m_{s}^{2}}+\frac{\sqrt{x}}{m_{i}m_{r}m_{s}}\Big)\\ &\ll\sum_{m_{s}\leq 2\sqrt{x+h}}\sum_{\frac{y}{m_{s}}<\tilde{m}\leq 2\sqrt{x+h}}3^{\nu(m_{s})}\Big(\frac{h}{\tilde{m}^{2}m_{s}^{2}}+\frac{\sqrt{x}}{\tilde{m}m_{s}}\Big)\\ &\ll\sum_{m_{s}\leq 2\sqrt{x+h}}3^{\nu(m_{s})}\Big(\frac{h}{ym_{s}}+\frac{\sqrt{x}}{m_{s}}\log x\Big)\\ &\ll\frac{h}{y}(\log x)^{3}+\sqrt{x}(\log x)^{4},\end{split}$
where the last bound follows from the elementary estimates $\sum_{m\leq x}\frac{1}{m}=\log x+O(1)$ and
@@ 5. Proof of Theorem 3
In [18, Sec. 4], the present author used the estimate
@@ Acknowledgments
The author would like to thank Professors Wen-Ching Winnie Li and Robert C. Vaughan for their suggestions. He is also very grateful to the referee for the careful reading and constructive comments.